OpenStudy (anonymous):

Solve: 4 = x+1 -- ----- x x+3

Parth (parthkohli):

\[\LARGE => 4(x + 3) = x(x + 1) \] \[\LARGE => 4x + 12 = x^{2} + 1x\] \[\LARGE => 3x = x^{2} - 12\] \[\LARGE => 3x + 12 = x^{2}\]